# Finite Element Operator Network for Solving Parametric PDEs

### v2

@@ 2.2 Convergence analysis of FEONet
We first note from (17) and (28), that $A$ defined in (13), (14) mostly determines the structures of the loss functions and it would be advantageous for us to analyze the loss function if we know more about the matrix. Note that the matrix $A$ is determined by the structure of the given differential equations, the choice of basis functions and the boundary conditions. Therefore, the characterization of $A$ which is useful for the analysis of the loss function and can cover a wide range of PDE settings simultaneously is important. The next lemma addresses this issue, which is quoted from [23].
@@
$\rho_{\min}|\bm{x}|\leq|A\bm{x}|\leq\rho_{\max}|\bm{x}|.$ (32)
Since the equation under consideration (9) is self-adjoint, the corresponding bilinear form $B[\cdot,\cdot]$ defined in (11) is symmetric which automatically guarantees that the matrix $A$ in our case is symmetric. Furthermore, due to the facts that the coefficient $\bm{a}(\cdot)$ is uniformly elliptic and $c(\cdot)$ is non-negative, the bilinear form $B[\cdot,\cdot]$ is coercive, which implies that $A$ is positive-definite. Therefore, we can apply Lemma 2.4 to the matrix $A$ of our interest.
@@
Assume that (10) holds. Then we have that
$\|\alpha^{*}-\widehat{\alpha}(n)\|_{L^{2}(\Omega)}\rightarrow 0\quad{\rm{as}}\,\,n\rightarrow\infty.$ (33)
@@
$\mathcal{G}_{n}:=\{|A\alpha-F|^{2}:\alpha\in\mathcal{N}_{n}\},$ (34)
where the matrix $A$ and the vector $F$ are defined in (15). In the following theorem, note that we assume that for any $n\in\mathbb{N}$, the Rademacher complexity of $\mathcal{G}_{n}$ converges to zero, which is known to be true in many cases [45, 14, 37]. For the proof of the theorem below, see Appendix A.2.
@@
Suppose that (10) holds, and we further assume that for any $n\in\mathbb{N}$, $\lim_{M\rightarrow\infty}R_{M}(\mathcal{G}_{n})=0$. Then we have with probability $1$ that
@@
$\lim_{n\rightarrow\infty}\lim_{M\rightarrow\infty}\|\alpha^{*}-\widehat{\alpha}(n,M)\|_{L^{2}(\Omega)}=0.$ (35)
Based on the convergence of the approximate coefficients (35), let us now state the convergence theorem for the predicted solution, whose proof is presented in Appendix A.3.
@@
Suppose (10) holds, and assume that for all $n\in\mathbb{N}$, $R_{M}(\widetilde{{\mathcal{G}}}_{n})\rightarrow 0$ as $M\rightarrow\infty$, with $\widetilde{{\mathcal{G}}}_{n}:=\{|A\alpha-F|^{2}:\alpha\in\mathcal{N}_{n}\}$. Then for given $h>0$, we have with probability $1$ over i.i.d. samples that
$\lim_{n\rightarrow\infty}\lim_{M\rightarrow\infty}\|u_{h}-u_{h,n,M}\|_{L^{2}(\Omega;L^{2}(D))}=0.$ (36)
@@ 3 Numerical experiments
In this section, we demonstrate the experimental results of the FEONet across various settings for PDE problems. We randomly generate 3000 input samples for external forces and train the FEONet in an unsupervised manner without the usage of pre-computed input-output $(f,u)$ pairs. We then evaluate the performance of our model with another randomly generated test set consisting of 3000 pairs of $f$ and the corresponding solutions $u$. For this test data, we employed the FEM with a sufficiently fine mesh discretization to ensure that the numerical solutions can be considered as true solutions (See Appendix B.1). While there are other operator-learning-based methods, such as PINO [30], that can be trained without input-output data pairs, they often face limitations when dealing with domains of complex shapes. Consequently, our primary comparison will be against the DeepONet [34], taking into account varying numbers of training data, and PIDeepONet [49]. The triangulation of domains and the generation of the nodal basis for the FEM are performed using FEniCS [2, 33]. A detailed description of the model used in the experiments can be found in Appendix B.
In Section 3.1, we will solve PDE problems in various environments and settings to verify the computational flexibility of our method. In Section 3.2, we will confirm the flexibility of the proposed operator learning method by employing various types of input functions, and in Section 3.3, we will demonstrate that the FEONet framework can make an accurate solution prediction for the singular perturbation problems. Finally, in Section 3.4, additional experiments will be conducted to confirm the robustness and reliability of our method across various performance metrics. Throughout the experiments, we will cover both one-dimensional and two-dimensional cases, and we have appropriately organized them according to the objectives of each experiment.
Table 1: Mean Rel. $L^{2}$ test errors (×$\mathbf{(\times 10^{-2})}$ with standard deviations (×$\mathbf{(\times 10^{-2})}$ for the 3000 test set on diverse PDE problems. Five training trials are performed independently. Domain I (circle), Domain II (square with a hole), Domain III (polygon) with Eq. (37), BC I (Dirichlet), BC II (Neumann) with Eq. (38), Eq I (second-order linear equation (39)) and Eq II (Burgers’ equation (40)).
@@
3.1 Performance evaluation in various scenarios
@@
$-\varepsilon\Delta u+\bm{v}\cdot\nabla u$ $=f(x,y),$ $(x,y)\in D,$ (37)
@@
Various boundary conditions
@@
$-\varepsilon u_{xx}+bu_{x}$ $=f(x),$ $x\in D,$ (38)
$u(x)=0\quad\text{or}\quad$ $u_{x}(x)=0,$ $x\in\partial D,$
where $\varepsilon=0.1$ and $b=-1$. For the Neumann boundary condition, we add $u$ to the left-hand side of (38) to address the uniqueness of the solution. The FEONet has an additional significant advantage to make the predicted solutions satisfy the exact boundary condition, by selecting the appropriate coefficients for the basis functions. Using the FEONet without any input-output training data, we are able to obtain similar or slightly higher accuracy compared to the DeepONet with 3000 training data, for both homogeneous Dirichlet or Neumann boundary conditions as shown in Table 1.
Various equations
@@
Since the Burgers’ equation is nonlinear, an iterative method is required for classical numerical schemes including the FEM, which usually causes some computational costs. Furthermore, it is more difficult than linear equations to obtain training data for nonlinear equations. One important advantage of the FEONet is its ability to effectively learn the nonlinear structure without any training data. Once the training process is completed, the model can provide accurate real-time solution predictions without relying on iterative schemes. The 7th and 8th columns of Table 1 highlight the effectiveness of the FEONet, predicting the solutions with reasonably low errors for various equations. Although the 8th column of Table 1 shows that the DeepONet with 3000 training data has a lower error compared to our model, it demonstrates the strength of the FEONet that achieves a similar level of accuracy even for the nonlinear PDE, that can be trained in an unsupervised manner.
Figure 4: Solution profiles predicted by the trained FEONet for PDE (39) obtained from variable coefficients $c(x)$ as an input function.
3.2 Various types of input functions
In this section, we will demonstrate that our method can accommodate various types of input functions. Specifically, we shall utilize, as an input for our model, variable coefficients, boundary conditions and initial conditions for time-dependent problems. Furthermore, for the case of the vector-valued input functions, we will also include the solution prediction for the 2D Stokes problem. Throughout the series of experiments, we will confirm that our approach can make a fast and accurate solution prediction for such various scenarios.
@@ 3.2 Various types of input functions
The FEONet is capable of learning the operator mapping $\mathcal{G}:c(x)\mapsto u(x)$ from the variable coefficient $c(x)$ to the corresponding solution $u(x)$ of a PDE (39). Figure 4 displays the solution profile obtained when variable coefficients are considered as input. This highlights the flexibility of FEONet in adapting to a wide range of input function types, thereby underscoring its versatility.
Figure 5: Solution profiles predicted by the trained FEONet for PDE (41) obtained from the boundary condition input. Note that the boundary condition $g(x,y)$ for the region with a circle hole inside $\Omega_{\text{cir}}$ is represented as $h(\theta)$ in polar coordinates. The average of relative $L^{2}$ errors across the entire test set when comparing the exact solution to the predictions is 0.0043.
@@
where $f(x,y)=2$ and the domain $D$ is a square with a hole (see (b) in Figure 1). The FEONet is utilized to learn the operator $\mathcal{G}:g(x,y)\mapsto u(x,y)$. Figure 5 demonstrates that even in the case of an operator that takes the boundary condition $g(x,y)$ as input, we can predict the solution with high accuracy. This demonstrates the versatility and adaptability of the FEONet approach for accommodating different types of input functions.
Initial condition as an input
Figure 6: Solution profiles predicted by the trained FEONet for PDE (42) obtained from the initial condition $u(t=0,x)$ as an input function. The average of relative $L^{2}$ error across the entire test set, when comparing the exact solution profile with the predicted solution, is 0.0010.
The FEONet can be extended to handle the initial conditions as input. We consider the time-dependent problem with varying initial conditions. We consider the 1D convection-diffusion equation
$u_{t}-\nu u_{xx}+bu_{x}$ $=0,$ $t\in[0,1],x\in D=[-1,1],$ (42)
@@
where $\nu=0.1$ and $b=-1$. We aim to learn the operator $\mathcal{G}:u_{0}(x)\mapsto u(t,x)$. We use the implicit Euler method to discretize the time with $\Delta t=0.01$. We employed the marching-in-time scheme proposed by [24] to sequentially learn and predict solutions over the time domain $t=[0,1]$, divided into 10 intervals. For training FEONet, we utilized only the input function $u_{0}(x)$ and equation (42) without any additional data. Figure 6 illustrates FEONet’s effective prediction of true solutions using an initial condition from the test dataset.
Vector-valued function as an input: 2D Stokes equation
Figure 7: Solution profiles predicted by the trained FEONet for 2D Stokes equation (43). The average of relative $L^{2}$ errors across the entire test set when comparing the exact solution to the predictions for $u_{1}$, $u_{2}$, and $p$ are 0.0499, 0.1092, and 0.0685, respectively.
@@
$-\nabla\cdot(\nabla\bm{u}-pI)$ $=\bm{f},$ $(x,y)\in D=[0,1]^{2},$ (43)
@@
where $\bm{g}(x,y)=(3+1.7\sin(2\pi x),0)$. We aim to learn the operator $\mathcal{G}:\bm{f}=(f_{1},f_{2})\mapsto[\bm{u},p]$ using the FEONet. For this, we employ the P2 Lagrange element for velocity and the P1 Lagrange element for pressure. Together, this combination is known as the Taylor-Hood element [15]. Figure 7 demonstrates the results of the operator learning for the 2D Stokes equation using FEONet. This illustrates that FEONet is applicable to the learning of operators for the system of PDEs with vectorial inputs and outputs.
@@ 3.3 Singular perturbation problem
One of the notable advantages of the FEONet, which predicts coefficients through well-defined basis functions, is its applicability to problems involving singularly perturbed PDEs. The boundary layer problem is a well-established challenge in scientific computing. This issue arises from the presence of thin regions near boundaries where the solution exhibits steep gradients. Such regions significantly impact the overall behavior of the system. Typically, in cases where the target function demonstrates abrupt transitions, such as when the diffusion coefficient is small in our model problem, neural network algorithms frequently encounter difficulties in converging to optimal solutions. This is largely due to the phenomenon known as spectral bias [43]. Neural networks are generally predicated on a smooth prior assumption, but spectral bias leads to a failure to accurately capture abrupt transitions or singular behaviors of the target solution function. To address this challenge, we propose a novel semi-analytic machine learning approach, guided by singular perturbation analysis, for effectively capturing the behavior of thin boundary layer. This is accomplished by incorporating additional basis functions guided by theoretical considerations. For this we consider
$-\varepsilon u_{xx}+bu_{x}$ $=f(x),$ $x\in D,$ (44)
@@
$\phi_{\text{cor}}(x):=e^{-(1+x)/\varepsilon}-(1-(1-e^{-2/\varepsilon})(x+1)/2).$ (45)
@@
By integrating the boundary layer element into the finite element space, we establish the proposed enriched Galerkin space for utilization in the FEONet. As the corrector basis is incorporated alongside the conventional nodal basis functions in the FEM, we also predict the additional coefficient originating from the corrector basis. Note that the computational cost remains minimal as the enriched basis exclusively encompasses boundary elements. Table 2 summarizes the characteristics of various operator learning models and the corresponding errors when solving the singularly perturbed PDE described in (44). The performance of the FEONet in accurately solving boundary layer problems surpasses that of other models, even without labeled data, across multiple instances. The incorporation of a very small $\varepsilon$ in the residual loss of the PDE makes operator learning challenging or unstable when using PINN and PIDeepONet. In fact, neural networks inherently possess a smooth prior, which can lead to difficulties in handling boundary layer problems. In contrast, FEONet utilizes theory-guided basis functions, enabling the predicted solution to accurately capture the sharp transition near the boundary layer. The PIDeepONet faces difficulties in effectively learning both the residual loss and boundary conditions, while PINNs struggle to predict the solution. Figure 8 exhibits the results of operator learning for the singular perturbation problem. As depicted in the zoomed-in graph of Figure 8, the FEONet stands out as the only model capable of accurately capturing sharp transitions, whereas other models exhibit noticeable errors.
3.4 Further performance metrics
Generalization errors

### v3

@@ 2.2 Convergence analysis of FEONet
We first note from (2.16) and (2.27), that $A$ defined in (2.12), (2.13) mostly determines the structures of the loss functions and it would be advantageous for us to analyze the loss function if we know more about the matrix. Note that the matrix $A$ is determined by the structure of the given differential equations, the choice of basis functions, and the boundary conditions. Therefore, the characterization of $A$ which is useful for the analysis of the loss function and can cover a wide range of PDE settings simultaneously is important. The next lemma addresses this issue, which is quoted from [26].
@@
$\rho_{\min}|\bm{x}|\leq|A\bm{x}|\leq\rho_{\max}|\bm{x}|.$ (2.31)
Since the equation under consideration (2.8) is self-adjoint, the corresponding bilinear form $B[\cdot,\cdot]$ defined in (2.10) is symmetric which automatically guarantees that the matrix $A$ in our case is symmetric. Furthermore, due to the facts that the coefficient $\bm{a}(\cdot)$ is uniformly elliptic and $c(\cdot)$ is non-negative, the bilinear form $B[\cdot,\cdot]$ is coercive, which implies that $A$ is positive-definite. Therefore, we can apply Lemma 2.5 to the matrix $A$ of our interest.
@@
Assume that (2.9) holds. Then we have that
$\|\alpha^{*}-\widehat{\alpha}(n)\|_{L^{2}(\Omega)}\rightarrow 0\quad{\rm{as}}\,\,n\rightarrow\infty.$ (2.32)
@@
$\mathcal{G}_{n}:=\{|A\alpha-F|^{2}:\alpha\in\mathcal{N}_{n}\},$ (2.33)
where the matrix $A$ and the vector $F$ are defined in (2.14). In the following theorem, note that we assume that for any $n\in\mathbb{N}$, the Rademacher complexity of $\mathcal{G}_{n}$ converges to zero, which is known to be true in many cases [48, 14, 40]. For the proof of the theorem below, see Appendix A.2.
@@
Suppose that (2.9) holds, and we further assume that for any $n\in\mathbb{N}$, $\lim_{M\rightarrow\infty}R_{M}(\mathcal{G}_{n})=0$. Then we have with probability $1$ that
@@
$\lim_{n\rightarrow\infty}\lim_{M\rightarrow\infty}\|\alpha^{*}-\widehat{\alpha}(n,M)\|_{L^{2}(\Omega)}=0.$ (2.34)
Based on the convergence of the approximate coefficients (2.34), let us now state the convergence theorem for the predicted solution, whose proof is presented in Appendix A.3.
@@
Suppose (2.9) holds, and assume that for all $n\in\mathbb{N}$, $R_{M}(\widetilde{{\mathcal{G}}}_{n})\rightarrow 0$ as $M\rightarrow\infty$, with $\widetilde{{\mathcal{G}}}_{n}:=\{|A\alpha-F|^{2}:\alpha\in\mathcal{N}_{n}\}$. Then for given $h>0$, we have with probability $1$ over i.i.d. samples that
$\lim_{n\rightarrow\infty}\lim_{M\rightarrow\infty}\|u_{h}-u_{h,n,M}\|_{L^{2}(\Omega;H^{s}(D))}=0.$ (2.35)
@@ 3 Numerical experiments
In this section, we demonstrate the experimental results of the FEONet across various settings for PDE problems. We randomly generate 3000 input samples for external forces and train the FEONet in an unsupervised manner without the usage of pre-computed input-output $(f,u)$ pairs. We then evaluate the performance of our model with another randomly generated test set consisting of 3000 pairs of $f$ and the corresponding solutions $u$. For this test data, we employed the FEM with a sufficiently fine mesh discretization to ensure that the numerical solutions can be considered as true solutions (See Appendix B.1). While there are other operator learning-based methods, such as PINO [33], that can be trained without input-output data pairs, they often face limitations when dealing with domains of complex shapes. Our primary comparison will be against DeepONet [37] and POD-DeepONet [38], with varying numbers of training data, as well as against PIDeepONet [53]. The FEONet is designed to learn the operator without requiring pre-computed training data, which significantly reduces computational overhead. The generation of random input samples, which are used for unsupervised learning, imposes negligible computational costs, offering flexibility in choosing the number of input function samples for training. In contrast, DeepONet (and POD-DeepONet) relies on supervised learning, requiring input-output data pairs. This necessitates pre-computing the output data for each input function sample, resulting in substantial computational demands. Consequently, it is challenging to directly compare the training times of FEONet with those of DeepONet (or POD-DeepONet). Note that training FEONet to solve the 2D Poisson equation in a domain with a circular hole, achieving a 1.3% relative error, took approximately 28 minutes and 36 seconds on an NVIDIA RTX A5000 24GB GPU. The triangulation of domains and the generation of the nodal basis for the FEM are performed using FEniCS [2, 36]. A detailed description of the model used in the experiments, as well as the comparison of computational time between FEM and operator learning models, can be found in Appendix B.2 and B.3.
In Section 3.1, we will solve PDE problems in various environments and settings to verify the computational flexibility of our method and confirm the flexibility of the proposed operator learning method by employing various types of input functions. In Section 3.2, we will conduct some further experiments on the parabolic problem and the Stokes problem. In Section 3.3, we will demonstrate that the FEONet framework can make an accurate solution prediction for the singular perturbation problems. Finally, in Section 3.4, additional experiments will be conducted to confirm the robustness and reliability of our method across various performance metrics. Throughout the experiments, we will cover both one-dimensional and two-dimensional cases, and we have appropriately organized them according to the objectives of each experiment.
Table 1: Mean Rel. $L^{2}$ test errors (×$\mathbf{(\times 10^{-2})}$ with standard deviations (×$\mathbf{(\times 10^{-2})}$ for the 3000 test set on diverse PDE problems. Five training trials are performed independently. Domain I (circle), Domain II (square with a hole), Domain III (polygon) with Eq. (3.1), BC I (Dirichlet), BC II (Neumann) with Eq. (3.2), Eq I (second-order linear equation (3.3)) and Eq II (Burgers’ equation (3.4)).
@@
Table 2: Mean Rel. $L^{2}$ test errors (×$\mathbf{(\times 10^{-2})}$ for the 1000 test set on diverse PDE problems. Coefficient (Variable coefficient as an input) with Eq. (3.3), Boundary (Boundary condition as an input) with Eq. (3.5), Initial (Time-dependent problem) with Eq. (3.6) and Vector-valued (Stokes problem) with Eq. (3.7).
Model (#Train data) Coefficient Boundary Initial Vector-valued
POD-DeepONet (supervised, w/10) 25.276 47.500 7446.077 -
POD-DeepONet (supervised, w/100) 10.536 28.471 230.288 -
POD-DeepONet (supervised, w/1000) 0.371 6.095 10.628 -
Ours (w/o labeled data) 0.612 0.430 0.218 5.141
3.1 Performance evaluation of FEONet in various scenarios
@@ 3.1.1 Simulations on different domains
$-\varepsilon\Delta u+\bm{v}\cdot\nabla u$ $=f(x,y),$ $(x,y)\in D,$ (3.1)
@@
3.1.2 Simulations on Dirichlet and Neumann boundary conditions
@@
$-\varepsilon u_{xx}+bu_{x}$ $=f(x),$ $x\in D,$ (3.2)
$u(x)=0\quad\text{or}$ $u_{x}(x)=0,$ $x\in\partial D,$
where $\varepsilon=0.1$ and $b=-1$. For the Neumann boundary condition, we add $u$ to the left-hand side of (3.2) to address the uniqueness of the solution. The FEONet has an additional significant advantage to make the predicted solutions satisfy the exact boundary condition, by selecting the appropriate coefficients for the basis functions. Using the FEONet without any input-output training data, we are able to obtain similar or slightly higher accuracy compared to the DeepONet with 3000 training data, for both homogeneous Dirichlet or Neumann boundary conditions as shown in fifth and sixth columns in Table 1.
3.1.3 Simulations on linear and nonlinear PDEs
@@
Since the Burgers’ equation is nonlinear, an iterative method is required for classical numerical schemes including the FEM, which usually causes some computational costs. Furthermore, it is more difficult than linear equations to obtain training data for nonlinear equations. One important advantage of the FEONet is its ability to effectively learn the nonlinear structure without any training data. Once the training process is completed, the model can provide accurate real-time solution predictions without relying on iterative schemes. The seventh and eighth columns of Table 1 highlight the effectiveness of the FEONet, predicting the solutions with reasonably low errors for various equations. Although the eighth column of Table 1 shows that the DeepONet with 3000 training data has a lower error compared to our model, it demonstrates the strength of the FEONet that achieves a similar level of accuracy even for the nonlinear PDE, that can be trained in an unsupervised manner.
Figure 4: Solution profiles predicted by the trained FEONet for PDE (3.3) obtained from variable coefficients $c(x)$ as an input function.
3.1.4 Simulations on various types of input functions
In this section, we will demonstrate that our method can accommodate various types of input functions. Specifically, we shall utilize, as an input for our model, variable coefficients and boundary conditions. Throughout the series of experiments, we will confirm that our approach can make a fast and accurate solution prediction for such various scenarios. Note, however, that the proposed method cannot take the geometry of the domain as an input of neural networks, which is of independent interest (See, for example, the references [54, 49] where this issue is addressed).
@@ 3.1.4 Simulations on various types of input functions
The FEONet is capable of learning the operator mapping $\mathcal{G}:c(x)\mapsto u(x)$ from the variable coefficient $c(x)$ to the corresponding solution $u(x)$ of a PDE (3.3). Figure 4 displays the solution profile obtained when variable coefficients are considered as input. The second column in Table 2 shows that when the input is a coefficient variable, POD-DeepONet, using all 1,000 data pairs, produces a slightly smaller error than FEONet, but the errors are quite similar. This demonstrates the strength of FEONet in achieving comparable accuracy while being trained in an unsupervised manner.
Figure 5: Solution profiles predicted by the trained FEONet for PDE (3.5) obtained from the boundary condition input. Note that the boundary condition $g(x,y)$ for the region with a circle hole inside $\Omega_{\text{cir}}$ is represented as $h(\theta)$ in polar coordinates. The average of relative $L^{2}$ errors across the entire test set when comparing the exact solution to the predictions is 0.0043.
@@
where $f(x,y)=2$ and the domain $D$ is a square with a hole (see (b) in Figure 1). The FEONet is utilized to learn the operator $\mathcal{G}:g(x,y)\mapsto u(x,y)$. Figure 5 demonstrates that even in the case of an operator that takes the boundary condition $g(x,y)$ as input, the FEONet can predict the solution with high accuracy, as shown in the third column of Table 2, compared to POD-DeepONet. This demonstrates the versatility and adaptability of the FEONet approach for accommodating different types of input functions.
3.2 Further extensions of FEONet
In this section, we will extend FEONet to handle the parabolic problem and the Stokes problem.
3.2.1 Time-dependent problem
The FEONet can be extended to handle the parabolic problem with initial conditions as input. We consider the time-dependent problem with varying initial conditions. We consider the 1D convection-diffusion equation
$u_{t}-\nu u_{xx}+bu_{x}$ $=0,$ $t\in[0,1],x\in D=[-1,1],$ (3.6)
@@ 3.2.1 Time-dependent problem
where $\nu=0.1$ and $b=-1$. We aim to learn the operator $\mathcal{G}:u_{0}(x)\mapsto u(t,x)$. We use the implicit Euler method to discretize the time with $\Delta t=0.01$. We employed the marching-in-time scheme proposed by [27] to sequentially learn and predict solutions over the time domain $t=[0,1]$, divided into 10 intervals. For training FEONet, we utilized only the input function $u_{0}(x)$ and equation (3.6) without any additional data. See Appendix B.4 for more details. Figure 6 illustrates FEONet’s effective prediction of true solutions using an initial condition from the test dataset. The fourth column in Table 2 shows that the FEONet—trained without any data pairs—yields a smaller error than POD-DeepONet.
Figure 6: Solution profiles predicted by the trained FEONet for PDE (3.6) obtained from the initial condition $u(t=0,x)$ as an input function. The average of relative $L^{2}$ error across the entire test set, when comparing the exact solution profile with the predicted solution, is 0.0022.
3.2.2 Stokes problem
Figure 7: Solution profiles predicted by the trained FEONet for 2D Stokes equation (3.7). The average of relative $L^{2}$ errors across the entire test set when comparing the exact solution to the predictions for $u_{1}$, $u_{2}$, and $p$ are 0.0499, 0.1092, and 0.0685, respectively.
@@ 3.2.2 Stokes problem
$-\nabla\cdot(\nabla\bm{u}-pI)$ $=\bm{f},$ $(x,y)\in D=[0,1]^{2},$ (3.7)
@@
where $\bm{g}(x,y)=(3+1.7\sin(2\pi x),0)$. We aim to learn the operator $\mathcal{G}:\bm{f}=(f_{1},f_{2})\mapsto[\bm{u},p]$ using the FEONet. For this, we employ the P2 Lagrange element for velocity and the P1 Lagrange element for pressure, which is known as the Taylor-Hood element [15]. Figure 7 demonstrates the results of the operator learning for the 2D Stokes equation using FEONet. This illustrates that FEONet is applicable to the learning of operators for the system of PDEs with vectorial inputs and outputs. In this problem, which requires predicting three outputs—$u_{1}(x,y)$, $u_{2}(x,y)$, and $p(x,y)$—the experiment for POD-DeepONet was omitted as shown in the fifth column of Table 2 because constructing a natural extension of the basis using POD is challenging. This remains an area for future work involving the application of POD-DeepONet.
@@ 3.3 Singular perturbation problem
Since the proposed method is based on domain meshing, we may face some potential challenges associated with the scenarios involving singular behavior including interior or boundary shocks. In particular, as our method is not mesh-free, refined meshes are required to handle certain problems, such as shock formation. This is a recognized limitation of our approach. However, while our method may face difficulties in solving singular problems, it benefits from the ability to leverage various existing numerical analysis techniques to address these challenges. To the best of our knowledge, such problems are very challenging for other machine learning-based approaches, including mesh-free methods. More precisely, one of the notable advantages of the FEONet, which predicts coefficients through well-defined basis functions, is its applicability to problems involving singularly perturbed PDEs. The boundary layer problem is a well-established challenge in scientific computing. This issue arises from the presence of thin regions near boundaries where the solution exhibits steep gradients. Such regions significantly impact the overall behavior of the system. Typically, in cases where the target function demonstrates abrupt transitions, such as when the diffusion coefficient is small in our model problem, neural network algorithms frequently encounter difficulties in converging to optimal solutions. This is largely due to the phenomenon known as spectral bias [46]. Neural networks are generally predicated on a smooth prior assumption, but spectral bias leads to a failure to accurately capture abrupt transitions or singular behaviors of the target solution function. To address this challenge, we propose a novel semi-analytic machine learning approach, guided by singular perturbation analysis, for effectively capturing the behavior of thin boundary layers. This is accomplished by incorporating additional basis functions guided by theoretical considerations. For this we consider
$-\varepsilon u_{xx}+bu_{x}$ $=f(x),$ $x\in D,$ (3.8)
@@
$\phi_{\text{cor}}(x):=e^{-(1+x)/\varepsilon}-(1-(1-e^{-2/\varepsilon})(x+1)/2).$ (3.9)
@@
By integrating the boundary layer element into the finite element space, we establish the proposed enriched Galerkin space for utilization in the FEONet. As the corrector basis is incorporated alongside the conventional nodal basis functions in the FEM, we also predict the additional coefficient originating from the corrector basis. Note that the computational cost remains minimal as the enriched basis exclusively encompasses boundary elements. Table 3 summarizes the characteristics of various operator learning models and the corresponding errors when solving the singularly perturbed PDE described in (3.8). The performance of the FEONet in accurately solving boundary layer problems surpasses that of other models, even without labeled data, across multiple instances. The incorporation of a very small $\varepsilon$ in the residual loss of the PDE makes operator learning challenging or unstable when using PINN and PIDeepONet. In fact, neural networks inherently possess a smooth prior, which can lead to difficulties in handling boundary layer problems. In contrast, FEONet utilizes theory-guided basis functions, enabling the predicted solution to accurately capture the sharp transition near the boundary layer. The PIDeepONet faces difficulties in effectively learning both the residual loss and boundary conditions, while PINNs struggle to predict the solution. Figure 8 exhibits the results of operator learning for the singular perturbation problem. As depicted in the zoomed-in graph of Figure 8, the FEONet stands out as the only model capable of accurately capturing sharp transitions, whereas other models exhibit noticeable errors.
3.4 Generalization capacity and convergence rate
Generalization capacity